Fermi National Laboratory
 

Operation Summary for September 9, 2011 - September 12, 2011

 

The Integrated Luminosity for the period from 9/5/11 to 9/12/11 was estimated at 47.62 pb-1 (inverse picobarns). NuMI reported receiving 8.74E18 protons on target during this same period.

Friday September 9

The day shift began with the Tevatron (TEV) conducting studies, with the Main Injector (MI) and Recycler off for an MI access to fill a kicker (KPS3A) with Fluorinert, with Operations monitoring the Antiproton Source (Pbar) stack (196.7E10), with the Linac and Booster operating normally, with the NuMI and MiniBooNE experiments off due to the MI access, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next TEV store.

During the day shift, the MI access ended at 9:32 AM. Operators put the TEV through a dry squeeze at 9:55 AM. Operations established beam to MiniBooNE at 10:23 AM. At 10:45 AM, operators reported a problem with the MI-40 upper bus power supply; operators contacted Power Supply personnel who reported that there were blown fuses on the switch gear in the yard. Power Supply personnel reported at 11:45 AM that they had replaced the fuses, but they also found a bad PLC, which they replaced. MI came back online at 1:31 PM and Pbar resumed stacking at 1:33 PM. TEV personnel began a low energy study around 1:45 PM. Operators put the TEV through a dry squeeze at 3:44 PM.

During the evening shift, Operations established a 450 GeV 3x3 study store around 4:30 PM. Operators reported at 5:17 PM that the NuMI beam permit tripped off due to bad beam positions; it reset. At 6:26 PM, the NuMI beam permit tripped off again on bad beam positions; operators lowered the MI beam intensity. Operations terminated the 3x3 study store and put the TEV through a dry squeeze at 11:42 PM.

Saturday September 10

The midnight shift began with Operations monitoring the stash (351.5E10), with Pbar stacking (28.8E10), and with the NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established a 3x3 study store at 1:03 AM and terminated the store at 3 AM. Operations put the TEV through a dry squeeze at 3:33 AM. Operations established a 3x3 study store at 4:39 AM and terminated the store at 7:17 AM. Operations put the TEV through a dry squeeze at 7:38 AM.

During the day shift, Operations established a 3x3 study store around 9:25 AM and terminated at 12:05 PM. Operations established a 3x3 study store around 1:37 PM and terminated at 4:02 PM.

During the evening shift, Operations put the TEV into HEP shot setup at 8:28 PM and established store 9097 at 9:40 PM with an average initial luminosity (AIL) of 360.71E30.

Sunday September 11

The midnight shift began with Operations monitoring store 9097 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators reported at 12:23 AM that the NuMI beam permit tripped off due to bad beam positions; it reset. At 7:20 AM, the NuMI beam permit tripped off again due bad beam positions; it reset.

During the day shift, Operations terminated store 9097 at 9:45 AM, put the TEV into shot setup, and established store 9098 at 11:09 AM with an AIL of 368.41E30. At 12:16 PM, the NuMI beam permit tripped off again due to bad beam positions; it reset. At 12:55 PM, a Pbar Accumulator separator (A:ISEP2V) tripped off; operators called in Separator personnel to investigate. The Separator personnel found and replaced a bad interlock module at 3:16 PM.

During the evening shift, at 5:09 PM, store 9097 aborted due to a bad CAMAC card, which was replaced. Operations put the TEV into shot setup at 5:54 PM and established store 9099 at 6:56 PM with an AIL of 238.65E30.

Monday September 12

The midnight shift began with Operations monitoring store 9099 and the stash, with Pbar stacking, and with the NuMI and MiniBooNE taking beam.

During the midnight shift, at 4:58 AM, operators reported that store 9099 aborted. Operators investigated and discovered a TEV sector C3 standby power unit failure; operators called in Power Supply personnel, who had the unit replaced by 6:13 AM. Operators began ramping the TEV at 7:04 AM. Operators reported that the MI-10 upper bus power supply tripped off on a bypass interlock loop trip; it reset. Operators reported that Cryo System personnel replaced a 5-volt power supply for the cryo system crate at F3. Operators reported that the NuMI beam permit tripped off at 7:16 AM; it reset. Operators put the TEV through a dry squeeze at 7:44 AM.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to conduct a study.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/12/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory